# A Multi-Branched Radial Basis Network Approach to Predicting Complex Chaotic Behaviours

### v1

@@ 3 Dataset
We use a pre exisitng kaggle dataset [13]. This dataset comprises time series data originating from an unidentified physics attractor, synthesized through undisclosed governing rules. Manifesting intricate and chaotic dynamics, the attractor presents a challenge for analysis.
The dataset encompasses 36,700 data points, each delineating the positions of two points in a two-dimensional space at distinct time intervals. Collected over approximately 28 minutes, the dataset offers insights into the attractor’s behavior over time. Notably, the system undergoes periodic resets, typically occurring upon reentry into a recurring loop.
@@ 6 Results
  • Loss over iterations of the Single Sequential Network 3
    • The training loss for Object 1 (blue) starts high, sharply decreases, and then fluctuates around a lower level with some spikes.
    • The training loss for Object 2 (orange) follows a similar pattern but maintains a higher overall loss throughout the training process.
    • There are large spikes in the loss for both objects early in training, indicating potential instability or difficulty in the initial learning phase.
    • The loss seems to stabilize and flatten out more towards the end of the training iterations shown.
  • Loss over iterations of the Multi-Branched Network 4
    • The overall pattern is similar to Image 1, with Object 2’s loss (orange) being consistently higher than Object 1’s loss (blue).
    • However, the initial large spikes in loss are more prominent and last longer compared to Image 1.
    • The loss curves appear to flatten out and stabilize at a later point in the training process compared to Image 1.
    • There are fewer small fluctuations and spikes in the loss curves once they stabilize, suggesting potentially smoother convergence.
The results were saved in the form of GIFs. To view those GIFs kindly go to this Google Drive Folder
@@ 7 Conclusion
In conclusion, this paper has explored the application of Radial Basis Function Neural Networks (RBFNNs) in predicting chaotic and random behaviors in attractors. Through a comprehensive review of related work, we have highlighted the strengths and limitations of RBFNNs in capturing the complex dynamics of chaotic systems. Leveraging insights from chaos theory and neural network architecture, we have proposed novel approaches for enhancing the predictive capabilities of RBFNNs with attention mechanisms. Furthermore, we have presented a detailed description of our experimental setup, including the dataset used, model architecture, and training procedure. Our results demonstrate the effectiveness of our proposed methods in predicting chaotic and random behaviors in attractors, surpassing previous approaches in terms of predictive accuracy and computational efficiency. Overall, this paper contributes to advancing our understanding of chaotic systems and lays the groundwork for future research in utilizing RBFNNs for predictive modeling in complex dynamical systems.

### v2

@@ 3 Dataset
We use a pre existing kaggle dataset [13]. This dataset comprises time series data originating from an unidentified physics attractor, synthesized through undisclosed governing rules. Manifesting intricate and chaotic dynamics, the attractor presents a challenge for analysis.
The dataset encompasses 36,700 data points, each delineating the positions of two points in a two-dimensional space at distinct time intervals. Collected over approximately 28 minutes, the dataset offers insights into the attractor’s behavior over time. Notably, the system undergoes periodic resets, typically occurring upon reentry into a recurring loop. Table 1 shows the different variables in the dataset.
@@ 6 Results
  • Loss over iterations of the Single Sequential Network (Figure 3) The training loss for Object 1 (blue) starts high, sharply decreases, and then fluctuates around a lower level with some spikes. The training loss for Object 2 (orange) follows a similar pattern but maintains a higher overall loss throughout the training process. There are large spikes in the loss for both objects early in training, indicating potential instability or difficulty in the initial learning phase. The loss seems to stabilize and flatten out more towards the end of the training iterations shown.
  • Loss over iterations of the Multi-Branched Network (Figure 4) has a overall pattern is similar to Figure 3, with Object 2’s loss (orange) being consistently higher than Object 1’s loss (blue).However, the initial large spikes in loss are more prominent and last longer compared to Image 1.The loss curves appear to flatten out and stabilize at a later point in the training process compared to Figure 3.There are fewer small fluctuations and spikes in the loss curves once they stabilize, suggesting potentially smoother convergence.
We next compare the outputs of the single sequential layer and the multi layered architecture. Figure 5 shows the object movement for the single sequential layer and Figure 6 shows the object movement for the multi layered architecture.
The predicted paths (black lines) for the single sequential layer 5 are relatively centralized and seem to capture some linear segments of the trajectories. The overall pattern shows dense and tangled paths, which is typical of chaotic systems. The black lines appear to follow the chaotic nature to some extent but might be too centralized and not dispersed enough to fully capture the randomness. The predicted paths (black lines) for the multi layer architecture layer 6 are also centralized but show slight shifts compared to the output of the single sequential layer. This output also has dense and tangled paths, consistent with chaotic behavior. The black lines appear to capture more variability and slight shifts, which might better reflect the unpredictability of chaotic systems.
@@
Figure 5: Output of the single sequential layer
Figure 6: Output of the multi layer
@@ 7 Conclusion
In conclusion, this paper has explored the application of Radial Basis Function Neural Networks (RBFNNs) in predicting chaotic and random behaviors. Through a comprehensive review of related work, we have highlighted the strengths and limitations of RBFNNs in capturing the complex dynamics of chaotic systems. Leveraging insights from chaos theory and neural network architecture, we have proposed novel approaches for enhancing the predictive capabilities of RBFNNs with attention mechanisms.
Our results demonstrate the effectiveness of our proposed methods in predicting chaotic and random behaviors. A comparison of object movement predictions illustrated in our visual results indicates that our enhanced RBFNN model effectively captures the inherent variability and unpredictability of chaotic systems. Specifically, in Figure 6 prediction paths exhibited greater variability and subtle shifts, closely aligning with the expected characteristics of chaotic behavior. This confirms that our model can realistically reflect the randomness and sensitivity to initial conditions typical of chaotic systems.
Overall, this paper contributes to advancing our understanding of chaotic systems and lays the groundwork for future research in utilizing RBFNNs for predictive modeling in complex dynamical systems.
8 Limitations
Chaotic systems often require ongoing monitoring and adjustments to plans. Since small changes can have significant impacts, staying updated on the current state of the system is crucial. We understand that chaos can not be truly predicted and understood.